# Renaissance: Investigating the Pretraining of Vision-Language Encoders

### v1

@@ 4.1 Premise
In our first set of experiments, we ask what is the effect of freezing the weights of various parts of the model during pretraining? Specifically, if we initialize the vision and text modules of a two-tower encoder with pretrained models from their respective domains, can we freeze one or both of these modules during pretraining? Freezing both modules means that we would only be pretraining the cross-modal and output layers of the model. Pretraining is usually the most compute intensive aspect of model development and we can reduce the GPU memory use and the overall compute required by freezing parts of the model. The compute savings would allow researchers to pretrain models that might otherwise be too large for their hardware requirements. Alternatively, they might also train smaller models at higher batch sizes and possibly obtain better results.
@@
Given that both vision and text encoder modules are pretrained in their respective domains, it makes intuitive sense that we might be able to skip at least some portion of their pretraining. These experiments should demonstrate empirically whether or not this is the case. Furthermore, the creators of the dual encoder [^1] Dual encoder models are a simpler model type that is currently not available on Renaissance. model LiT Zhai et al. (2022) found that they obtained slightly better results from freezing the model’s vision encoder. This experiment will also afford us the opportunity to see if a similar effect will hold for two-tower encoder models.
Text Encoder Vision Encoder SNLI-VE NLVR2 Ref. Res.
Unfrozen Unfrozen 0.741 0.672 0.724
Frozen Unfrozen 0.735 0.675 0.702
Unfrozen Frozen 0.741 0.672 0.740
Frozen Frozen 0.738 0.665 0.721
ELECTRA-Base-Frz ViT-Base-Frz 0.756 0.630 0.756
@@
All models are trained for 100k steps at a batch size of 704. All results are calculated on dev sets for each task.
4.2 Experimental Setup and Procedure
To begin, we use Renaissance to construct a set of two-tower models with ELECTRA-Small Clark et al. (2020) as the text encoder and DeiT-Tiny Touvron et al. (2021) as the image encoder. Both of these models are both quite small and efficient, and we chose them to expedite the training process. We set the cross-modal encoder module of each model to contain two sets of six transformer layers each with a hidden size of 256 and 4 attention heads. In total we pretrain four model variations, a baseline with both modules unfrozen, one with the text encoder frozen, one with the image encoder frozen and one with both encoder modules frozen. Each model will be pretrained for 100k steps at a batch size of 704 using the masked language modeling and image-text matching tasks described in Section 3.1.3. All models are trained using two NVIDIA L40s GPUs. We use two of the four pretraining datasets, MSCOCO and Visual Genome, which were described in the same section. Finally, we finetune and evaluate our models on three of the five VL tasks described in Section 3.1.4: SNLI-VE, NLVR2 and reference resolution with RefCOCO. In the interest of saving time and compute, we forego evaluating them on multimodal retrieval tasks and visual question answering. A crucial point to consider is that the no model weights are frozen during finetuning.
4.3 Results
The results for this experiment are summarized in Table 1. We see that we can obtain similar results by freezing one or both of the previously trained encoder modules during pretraining with only mild ill effect. On the SNLI-VE task, the difference between training the whole model and freezing one or both modules is very slight indeed. When freezing the vision module, the downstream results for SNLI-VE are essentially identical and we see only a slight drop in performance compared to the baseline model. We see a slightly different pattern on the NLVR2 task, however. Here we see the results for the baseline model and two models with a single encoder frozen having almost identical results. The model with the visual encoder produces the best score on the reference resolution task. The baseline model and the model with both modules frozen preform very nearly identically, while the the model with the text encoder only frozen scores the worst. Of the four models, the overall best performance is achieved by freezing only the vision encoder.
This is a fairly remarkable result as we can essentially cut the GPU memory required during pretraining in half by freezing so many of the model’s weights. This is especially significant because pretraining the model is by far the largest compute cost we see during training. We should also note that this effect is somewhat similar to a phenomenon noted in the training of the dual encoder LiT Zhai et al. (2022) (dual encoders have two encoder stacks but lack the a cross-modal fusion module). Here Zhai et al. found that they can obtain better results from freezing the image encoder during training. Though our model architecture is different, we observe a somewhat similar effect.
To further demonstrate the utility of freezing the pretrained modules, we train a model that uses ELECTRA-Base as a text-encoder and ViT-Base as the vision encoder. We use the same training hyperparameters but increase the number of cross-attention layers to 10. This model is quite large for an encoder, containing over 210M parameters (decoder models for generative tasks often contain billions of parameters). A model of this size would be well outside our compute capacity to pretrain if the text and vision modules were not frozen. However, because it contains less than 27M trainable parameters, training it puts the same memory load on our two L40s GPUs as the baseline model does. The results for this model are displayed in the final row of Table 1. This model obtains the best results of the study on two of the three downstream tasks and remains well within our limited compute budget.
Encoder SNLI-VE NLVR2 Ref. Res.
Random 0.699 0.551 0.554
ViT 0.685 0.534 0.522
BERT 0.692 0.545 0.507
@@ 4.3 Results
5 Experiment 2: Text Encoder vs. Vision Encoder
5.1 Premise
In the previous set of experiments we focused on training two-tower models. In our final experiment we will examine the behavior of one-tower models. One-tower encoder models were also described in Section 3.1.1. To date, most of these models have been derived from text encoder models such as BERT Devlin et al. (2019). A less explored approach is to base such models on transformer based vision models such as ViT Dosovitskiy et al. (2020); this is the approach of the one-tower VL transformer called ViLT Kim et al. (2021). In this experiment we ask if one strategy is superior to another when training and evaluating under otherwise similar conditions? More simply put, are one-tower encoders more effective when based on a vision encoder or a text encoder. In addition to providing guidance to future practitioners of VL modeling, answering this question should also provide interesting results from purely theoretical as well as practical perspectives.
5.2 Experimental Setup and Procedure
To make this experiment as fair a comparison as possible we select a vision transformer and a text transformer model as close in size to each other and architecture as possible. Toward this end we used BERT Devlin et al. (2019) as our text-encoder model and ViT Dosovitskiy et al. (2020) as our vision encoder model. The encoder towers in each of these models were consciously designed to have nearly identical dimensions with each encoder module containing 110M parameters. We employ patch embeddings for visual tokens and word-piece embeddings for visual tokens in all models. The resultant models will be close to identical, save that the weights of one are derived from vision pretraining and the other from text/language pretraining. As a baseline, we also train a randomly initialized version based on the BERT architecture. Finally, we train each model for 50k steps with a batch size of 512 using masked-language modeling and image-text matching. Again we use MSCOCO and Visual Genome as training datasets and evaluate on the 3 vision-language tasks described in the previous experiments.
5.3 Results
The results for this experiment are displayed in Table 2. According to our analysis there doesn’t appear to be a significant advantage in basing a one-tower encoder model on either text or vision. Surprisingly, the randomly initialized model that we trained as a baseline scored the best on all three downstream tasks. These are very much unexpected results. Though we didn’t have an intuition as to whether text or vision would perform better, we didn’t expect the downstream results to be so similar and to be inferior to a randomly initialized variation. These results are especially notable since one of the few in depth analyses of vision-language models, Frank et al. (2021), indicates that the interaction between the visual and language modalities are not symmetric. That study used probing techniques to show that VL transformers learn to use vision-for-language more than language-for-vision. Our best explanation of this phenomenon is that that one-tower models do not make use of the individual visual or textual modalities, but instead converge to values not dependent on either.
Another notable conclusion of this experiment and the preceding ones, is that two-tower models are in general much more parameter efficient than one-tower models. The one-tower models used in this experiment are relatively large, each containing more than 100M parameters. While the two-tower models in the previous experiments contain less than 40M parameters. Nonetheless, the two-tower models outperform those in this final experiment using the same datasets for training and evaluation. In previous studies, one-tower models such as ViLT Kim et al. (2021) that have similar architectures, have obtained better results than those displayed here. They do so by using more data and enormous pretraining batch sizes that require significantly more compute than we used here. Though only a preliminary finding, this insight might prove valuable to those interested in efficient VL modeling.
6 Future Directions
6.1 Renaissance
As this program evolves, we hope to incorporate a number of additional features that are not available in the current version. The capabilities that we plan to add are discussed below.
6.1.1 Model Types
There are several model types, beyond one-tower and two-tower encoders, that we hope to support in future versions. These include, dual encoder, encoder-decoder and decoder only model types (see Fields and Kennington (2023) for explanations of and examples for each type). By virtue of adding these model types, we also hope to include the ability to generate text for tasks such as image captioning.
6.1.2 Additional Tasks
@@
6.2 Analysis of VL Transformers and Pretraining
Because of the field of vision-language is rapidly evolving there are many possible future directions for research. We will mention a few. Though we have touched on some of the more basic aspects of training in this study, a systematic study of how each pretraining task contributes to downstream performance would be very illuminating. As would testing other tasks, such as visual grounding or visual question answering in pretraining. A thorough investigation of which architectures are best used in which circumstances would also be a worthwhile endeavor. As a final suggestion, we believe the field would also benefit from a scaling study to determine the optimum amount of data to train models at various scales such as Kaplan et al. (2020) performed for NLP transformers.
7 Conclusion
In this study, we have examined some basic features of pretraining vision-language transformers. In addition to the experiments that we’ve performed, we also introduced a flexible vision-language modeling framework called Renaissance, the source code for which can be found at https://github.com/bsu-slim/renaissance. In our first set of experiments we showed that pretrained vision and text modules can be frozen during vision-language pretraining with only small losses in downstream performance. This finding opens the possibility of training VL models whose size might normally exceed one’s compute budget. In our second and final experiment we compared of effect of basing a one-tower encoder model on a text transformer versus a vision transformer. Surprisingly, our results indicate that neither strategy is superior to the other and that randomly initializing model weights yields the best results. We therefore recommend training one-tower models from scratch when possible. We conclude this study with the observation that multi-modal modeling is a rapidly expanding pursuit and we hope that this paper is among the first of many that aim to shed light on this dynamic and exciting field of deep learning.
Limitations
The primary limitations of our study relate to size and scope. With greater resources, particularly compute resources, we would have been able to test more models on a wider variety of downstream tasks. These additional data would have added greater weight to our findings and given them broader applicability.
@@ 7 Conclusion
We have no pertinent ethical conflicts to report.
Acknowledgements
The code for Renaissance borrows large portions from the code bases of METER Dou et al. (2022) and ViLT Kim et al. (2021). Without these excellent templates this program could not have been completed on a reasonable time scale. We would also to like to acknowledge the programmers at Huggingface. Our programs design interfaces with classes from the hub and our design borrowed heavily from their methods.
Appendix A Hardware
In all of the three studies we pertrain all of our models using two NVIDIA L40S GPUs each with 48GB of GPU memory. Where feasible we also used a server with two NVIDIA TITAN RTX GPUs with 24GB of memory and a server with two NVIDA TITAN Xp servers with 12GB of memory for finetuning.

### v2

@@ 4.1.  Purpose
In our first set of experiments, we ask what is the effect of freezing the weights of various parts of the model during pretraining? Specifically, if we initialize the vision and text modules of a two-tower encoder with pretrained models from their respective domains, can we freeze one or both of these modules during pretraining? Freezing both modules means that we would only be pretraining the cross-modal and output layers of the model. Pretraining is usually the most compute intensive aspect of model development and we can reduce the GPU memory use and the overall compute required by freezing parts of the model. The compute savings would allow researchers to pretrain models that might otherwise be too large for their hardware to support. [^1] Alternatively, they might also train smaller models at higher batch sizes and possibly obtain better results, as was found in Fields and Kennington (2023a).
@@
Given that both vision and text encoder modules are pretrained in their respective domains, it makes intuitive sense that we might be able to skip at least some portion of their pretraining. These experiments should demonstrate empirically whether or not this is the case. Furthermore, the creators of the dual encoder [^2] Dual encoder models are a simpler model type that is currently not available on Renaissance. model LiT (Zhai et al., 2022) found that they obtained slightly better results from freezing the model’s vision encoder. This experiment will also afford us the opportunity to see if a similar effect will hold for two-tower encoder models.
4.2.  Experimental Setup and Procedure
To begin, we use Renaissance to construct a set of two-tower models with ELECTRA-Small (Clark et al., 2020) as the text encoder and DeiT-Tiny (Touvron et al., 2021) as the image encoder. Both of these models are both quite small and efficient, and we chose them to expedite the training process. We set the cross-modal encoder module of each model to contain two sets of six transformer layers each with a hidden size of 256 and 4 attention heads. In total, we pretrain four model variations, a baseline with both modules unfrozen, one with the text encoder frozen, one with the image encoder frozen and one with both encoder modules frozen. Each model is pretrained for 100k steps at a batch size of 704 using the masked language modeling and image-text matching tasks described in Section 3.1.3. All models are trained using two NVIDIA L40s GPUs. We use two of the four pretraining datasets, MSCOCO and Visual Genome, which were described in the same section. Finally, we finetune and evaluate our models on three of the five VL tasks described in Section 3.1.4: SNLI-VE, NLVR2 and reference resolution with RefCOCO. A crucial point to consider is that the no model weights are frozen during finetuning.
4.3.  Results
The results for this experiment are summarized in Table 1. We see that we can obtain similar results by freezing one or both of the previously trained encoder modules during pretraining with only mild ill effect. On the SNLI-VE task, the difference between training the whole model and freezing one or both modules is very slight indeed. When freezing the vision module, the downstream results for SNLI-VE are essentially identical and we see only a slight drop in performance compared to the baseline model. We see a slightly different pattern on the NLVR2 task, however. Here,the results for the baseline model and two models with a single encoder frozen have almost identical results. The model with the visual encoder frozen produces the best score on the reference resolution task. The baseline model and the model with both modules frozen preform very nearly identically, while the the model with the text encoder only frozen scores the worst. Of the four models, the overall best performance is achieved by freezing only the vision encoder.
This is a fairly remarkable result as we can significantly cut the GPU memory required during pretraining by freezing so many of the model’s weights. This is especially significant because pretraining the model is by far the largest compute cost we see during training. We should also note that this effect is somewhat similar to a phenomenon noted in the training of the dual encoder LiT (Zhai et al., 2022) (dual encoders have two encoder stacks but lack the a cross-modal fusion module). Here Zhai et al. found that they can obtain better results from freezing the image encoder during training. Though our model architecture is different, we observe a somewhat similar effect.
Text Encoder Vision Encoder SNLI-VE NLVR2 Ref. Res. Avg.
Unfrozen Unfrozen 0.741 0.672 0.724 0.712
Frozen Unfrozen 0.735 0.675 0.702 0.704
Unfrozen Frozen 0.741 0.672 0.740 0.717
Frozen Frozen 0.738 0.665 0.721 0.708
@@ 4.3.  Results
All models are trained for 100k steps at a batch size of 704, the largest size our server’s could accomodate. All results are calculated on dev sets for each task.
5.  Experiment 2: Text Encoder vs. Vision Encoder
5.1.  Purpose
In the previous set of experiments we focused on training two-tower models. In our final experiment, we examine the behavior of one-tower models (described in Section 3.1.1). To date, most of these models have been derived from text encoder models such as BERT (Devlin et al., 2019). A less explored approach is to base such models on transformer based vision models such as ViT (Dosovitskiy et al., 2020); this is the approach of the one-tower VL transformer called ViLT (Kim et al., 2021). In this experiment we ask if one strategy is superior to another when training and evaluating under otherwise similar conditions? More simply put, are one-tower encoders more effective when based on a vision encoder or a text encoder. In addition to providing guidance to future practitioners of VL modeling, answering this question should also provide interesting results from purely theoretical as well as practical perspectives.
5.2.  Experimental Setup and Procedure
To make this experiment as fair a comparison as possible, we select a vision transformer and a text transformer model as close in size to each other and architecture as possible. Toward this end we used BERT (Devlin et al., 2019) as our text-encoder model and ViT (Dosovitskiy et al., 2020) as our vision encoder model. The encoder towers in each of these models were consciously designed to have nearly identical dimensions with each encoder module containing 110M parameters. We employ patch embeddings for visual tokens and word-piece embeddings for visual tokens in all models. The resultant models will be close to identical, save that the weights of one are derived from vision pretraining and the other from text/language pretraining. As a baseline, we also train a randomly initialized version based on the BERT architecture. Finally, we train each model for 50k steps with a batch size of 512 using masked-language modeling and image-text matching. Again we use MSCOCO and Visual Genome as training datasets and evaluate on the 3 VL tasks described in the previous experiments.
5.3.  Results
The results for this experiment are displayed in Table 2. According to our analysis, there doesn’t appear to be a significant advantage in basing a one-tower encoder model on either text or vision. Surprisingly, the randomly initialized model that we trained as a baseline scored the best on all three downstream tasks. These are very much unexpected results. Though we didn’t have an intuition as to whether text or vision would perform better, we didn’t expect the downstream results to be so similar and to be inferior to a randomly initialized variation. These results are especially notable since one of the few in depth analyses of VL models, Frank et al. (2021), indicates that the interaction between the visual and language modalities are not symmetric. That study used probing techniques to show that VL transformers learn to use vision-for-language more than language-for-vision. Our best explanation of this phenomenon is that that one-tower models do not make use of the individual visual or textual modalities, but instead converge to values not dependent on either.
Another notable conclusion of this experiment and the preceding ones, is that two-tower models are in general much more parameter efficient than one-tower models. The one-tower models used in this experiment are relatively large, each containing more than 100M parameters. While the two-tower models in the previous experiments contain less than 40M parameters. Nonetheless, the two-tower models outperform those in this final experiment using the same datasets for training and evaluation. In previous studies, one-tower models such as ViLT (Kim et al., 2021) that have similar architectures, have obtained significantly better results than those displayed here. They do so by using more data and enormous pretraining batch sizes that require significantly more compute than we used here. Though only a preliminary finding, this insight might prove valuable to those interested in efficient VL modeling.
Encoder SNLI-VE NLVR2 Ref. Res. Avg.
Random 0.699 0.551 0.554 0.601
ViT 0.685 0.534 0.522 0.581
BERT 0.692 0.545 0.507 0.581
@@ 5.3.  Results
6.  Future Directions
As Renaissance evolves, we hope to incorporate a number of additional features that are not available in the current version. The capabilities that we plan to add are discussed below.
Model Types
There are several model types, beyond one-tower and two-tower encoders, that we plan to support in future versions. These include, dual encoder, encoder-decoder and decoder only model types (see Fields and Kennington (2023b) for explanations of and examples for each type). By virtue of adding these model types, we also hope to include the ability to generate text for tasks such as image captioning.
Additional Tasks
@@ 6.  Future Directions
Analysis of VL Transformers and Pretraining
Because of the field of VL is rapidly evolving there are many possible future directions for research. We will mention a few. Though we have touched on some of the more basic aspects of training in this study, a systematic study of how each pretraining task contributes to downstream performance would be very illuminating, as would testing other tasks such as visual grounding or visual question answering in pretraining. A thorough investigation of which architectures are best used in which circumstances would also be a worthwhile endeavor.
7.  Conclusion
In this study, we have examined some basic features of pretraining vision-language transformers. In addition to the experiments that we’ve performed, we also introduced a flexible vision-language modeling framework called Renaissance, the source code for which will by made available upon publication. In our first set of experiments we showed that pretrained vision and text modules can be frozen during vision-language pretraining with only small losses in downstream performance. This finding opens the possibility of training VL models whose size might normally exceed one’s compute budget. In our second and final experiment we compared of effect of basing a one-tower encoder model on a text transformer versus a vision transformer. Surprisingly, our results indicate that neither strategy is superior to the other and that randomly initializing model weights yields the best results. We therefore recommend training one-tower models from scratch when possible. We conclude this study with the observation that multi-modal modeling is a rapidly expanding pursuit and we hope that this paper is among the first of many that aim to shed light on this dynamic and exciting field of deep learning.
@@ 7.  Conclusion
The authors have no ethical conflicts to report.
8.  Bibliographical References
  • A. Aghajanyan, L. Yu, A. Conneau, W. Hsu, K. Hambardzumyan, S. Zhang, S. Roller, N. Goyal, O. Levy, and L. Zettlemoyer (2023) Scaling laws for generative mixed-modal language models. In International Conference on Machine Learning, pp. 265–279. Cited by: §2.1.
  • S. Antol, A. Agrawal, J. Lu, M. Mitchell, D. Batra, C. L. Zitnick, and D. Parikh (2015) Vqa: visual question answering. In Proceedings of the IEEE international conference on computer vision, pp. 2425–2433. Cited by: item 5.
  • H. Bao, L. Dong, S. Piao, and F. Wei (2021) Beit: bert pre-training of image transformers. arXiv preprint arXiv:2106.08254. Cited by: §3.1.1.
  • E. Bugliarello, R. Cotterell, N. Okazaki, and D. Elliott (2021) Multimodal pretraining unmasked: a meta-analysis and a unified framework of vision-and-language berts. Transactions of the Association for Computational Linguistics 9, pp. 978–994. Cited by: §1, §2.1.
  • E. Bugliarello, L. Sartran, A. Agrawal, L. A. Hendricks, and A. Nematzadeh (2023) Measuring progress in fine-grained vision-and-language understanding. arXiv preprint arXiv:2305.07558. Cited by: §2.1.
  • M. Caron, H. Touvron, I. Misra, H. Jégou, J. Mairal, P. Bojanowski, and A. Joulin (2021) Emerging properties in self-supervised vision transformers. In Proceedings of the IEEE/CVF international conference on computer vision, pp. 9650–9660. Cited by: §3.1.1.
  • Y. Chen, L. Li, L. Yu, A. El Kholy, F. Ahmed, Z. Gan, Y. Cheng, and J. Liu (2020) Uniter: universal image-text representation learning. In European conference on computer vision, pp. 104–120. Cited by: §3.1.1.
  • K. Clark, M. Luong, Q. V. Le, and C. D. Manning (2020) Electra: pre-training text encoders as discriminators rather than generators. arXiv preprint arXiv:2003.10555. Cited by: §3.1.1, §4.2.
  • J. Devlin, M. Chang, K. Lee, and K. Toutanova (2019) BERT: pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies, Volume 1 (Long and Short Papers), pp. 4171–4186. External Links: Document, Link Cited by: 1st item, §3.1.1, §3.1.1, §3.1.2, §5.1, §5.2.
  • A. Dosovitskiy, L. Beyer, A. Kolesnikov, D. Weissenborn, X. Zhai, T. Unterthiner, M. Dehghani, M. Minderer, G. Heigold, S. Gelly, et al. (2020) An image is worth 16x16 words: transformers for image recognition at scale. arXiv preprint arXiv:2010.11929. Cited by: §3.1.1, §3.1.1, §3.1.1, §3.1.2, §5.1, §5.2.
  • Z. Dou, Y. Xu, Z. Gan, J. Wang, S. Wang, L. Wang, C. Zhu, P. Zhang, L. Yuan, N. Peng, et al. (2022) An empirical study of training end-to-end vision-and-language transformers. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition, pp. 18166–18176. Cited by: §3.1.1.
  • C. Fields and C. Kennington (2023a) Exploring transformers as compact, data-efficient language models. In Proceedings of the 27th Conference on Computational Natural Language Learning, Cited by: footnote 1.
  • C. Fields and C. Kennington (2023b) Vision language transformers: a survey. arXiv preprint arXiv:2307.03254. Cited by: §1, §3.1.3, §6.
  • S. Frank, E. Bugliarello, and D. Elliott (2021) Vision-and-language or vision-for-language? on cross-modal influence in multimodal transformers. arXiv preprint arXiv:2109.04448. Cited by: §2.1, §5.3.
  • K. He, X. Zhang, S. Ren, and J. Sun (2015) Deep residual learning for image recognition. External Links: 1512.03385, Link Cited by: §3.1.1, §3.1.1.
  • Z. Huang, Z. Zeng, B. Liu, D. Fu, and J. Fu (2020) Pixel-bert: aligning image pixels with text by deep multi-modal transformers. arXiv preprint arXiv:2004.00849. Cited by: §3.1.1.
  • S. Kazemzadeh, V. Ordonez, M. Matten, and T. Berg (2014) Referitgame: referring to objects in photographs of natural scenes. In Proceedings of the 2014 conference on empirical methods in natural language processing (EMNLP), pp. 787–798. Cited by: item 3.
  • W. Kim, B. Son, and I. Kim (2021) Vilt: vision-and-language transformer without convolution or region supervision. In International Conference on Machine Learning, pp. 5583–5594. Cited by: §5.1, §5.3.
  • R. Krishna, Y. Zhu, O. Groth, J. Johnson, K. Hata, J. Kravitz, S. Chen, Y. Kalantidis, L. Li, D. A. Shamma, et al. (2017) Visual genome: connecting language and vision using crowdsourced dense image annotations. International journal of computer vision 123 (1), pp. 32–73. Cited by: §3.1.3.
  • D. Li, J. Li, H. Le, G. Wang, S. Savarese, and S. C.H. Hoi (2023) LAVIS: a one-stop library for language-vision intelligence. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 3: System Demonstrations), Toronto, Canada, pp. 31–41. External Links: Link Cited by: §2.2.
  • L. H. Li, M. Yatskar, D. Yin, C. Hsieh, and K. Chang (2019) Visualbert: a simple and performant baseline for vision and language. arXiv preprint arXiv:1908.03557. Cited by: §3.1.1.
  • T. Lin, M. Maire, S. Belongie, J. Hays, P. Perona, D. Ramanan, P. Dollár, and C. L. Zitnick (2014) Microsoft coco: common objects in context. In Computer Vision–ECCV 2014: 13th European Conference, Zurich, Switzerland, September 6-12, 2014, Proceedings, Part V 13, pp. 740–755. Cited by: item 4, §3.1.3.
  • Z. Liu, H. Mao, C. Wu, C. Feichtenhofer, T. Darrell, and S. Xie (2022) A convnet for the 2020s. External Links: 2201.03545, Link Cited by: §3.1.1, §3.1.1.
  • J. Lu, D. Batra, D. Parikh, and S. Lee (2019) Vilbert: pretraining task-agnostic visiolinguistic representations for vision-and-language tasks. Advances in neural information processing systems 32. Cited by: §3.1.1.
  • B. McKinzie, Z. Gan, J. Fauconnier, S. Dodge, B. Zhang, P. Dufter, D. Shah, X. Du, F. Peng, A. Belyi, et al. (2024) Mm1: methods, analysis and insights from multimodal llm pre-training. In European Conference on Computer Vision, pp. 304–323. Cited by: §2.1.
  • V. Ordonez, G. Kulkarni, and T. Berg (2011) Im2text: describing images using 1 million captioned photographs. Advances in neural information processing systems 24. Cited by: §3.1.3.
  • A. Rogers, O. Kovaleva, and A. Rumshisky (2021) A primer in bertology: what we know about how bert works. Transactions of the Association for Computational Linguistics 8, pp. 842–866. Cited by: §1.
  • P. Sharma, N. Ding, S. Goodman, and R. Soricut (2018) Conceptual captions: a cleaned, hypernymed, image alt-text dataset for automatic image captioning. In Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), Melbourne, Australia, pp. 2556–2565. External Links: Link, Document Cited by: §3.1.3.
  • A. Suhr, S. Zhou, A. Zhang, I. Zhang, H. Bai, and Y. Artzi (2018) A corpus for reasoning about natural language grounded in photographs. arXiv preprint arXiv:1811.00491. Cited by: §1, Figure 3, item 1.
  • H. Tan and M. Bansal (2019) Lxmert: learning cross-modality encoder representations from transformers. arXiv preprint arXiv:1908.07490. Cited by: §3.1.1.
  • P. Tong, E. Brown, P. Wu, S. Woo, A. J. V. IYER, S. C. Akula, S. Yang, J. Yang, M. Middepogu, Z. Wang, et al. (2024) Cambrian-1: a fully open, vision-centric exploration of multimodal llms. Advances in Neural Information Processing Systems 37, pp. 87310–87356. Cited by: §2.1.
  • H. Touvron, M. Cord, M. Douze, F. Massa, A. Sablayrolles, and H. Jégou (2021) Training data-efficient image transformers & distillation through attention. In International conference on machine learning, pp. 10347–10357. Cited by: §3.1.1, §4.2.
  • A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, Ł. Kaiser, and I. Polosukhin (2017) Attention is all you need. Advances in neural information processing systems 30. Cited by: §3.1.1.
  • N. Xie, F. Lai, D. Doran, and A. Kadav (2019) Visual entailment: a novel task for fine-grained image understanding. arXiv preprint arXiv:1901.06706. Cited by: item 2.
  • X. Xu, C. Wu, S. Rosenman, V. Lal, W. Che, and N. Duan (2023) Bridgetower: building bridges between encoders in vision-language representation learning. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 37, pp. 10637–10647. Cited by: §3.1.1.
  • X. Zhai, X. Wang, B. Mustafa, A. Steiner, D. Keysers, A. Kolesnikov, and L. Beyer (2022) Lit: zero-shot transfer with locked-image text tuning. In Proceedings of the IEEE/CVF conference on computer vision and pattern recognition, pp. 18123–18133. Cited by: §4.1, §4.3.